# Stabilizer Entanglement as a Magic Highway

### v2

@@ Appendix D Magic injection with shallow-depth brickwork circuits
Figure 7: Numerical results for $\overline{Y}$ under brickwork circuits of varying depths and gate ranges, as depicted in Fig. 6.
Appendix E Exact results
@@
In this section, we present an alternative analytical method to derive exact expressions of $\overline{Y}$ in all cases discussed in the main text [^1] We acknowledge Alioscia Hamma for providing insightful notes, which served as the primary inspiration for the analytical method presented in this section.. Recall the linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=D^{-1}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$. where $D:={\rm dim}\ \mathcal{H}$. This quantity admits an equivalent expression:
$Y(\rho)=D{\rm tr}(Q\rho^{\otimes 4}),$ (70)
@@
where $Q:=D^{-2}\sum_{P\in P_{N}}P^{\otimes 4}$. For a Hilbert space $\mathcal{H}=\otimes^{n}_{i=1}\mathcal{H}_{i}$, the operator $Q$ factories as $Q=\otimes^{n}_{i=1}Q_{i}$. We can then interpret the map $M(\cdot):=D{\rm tr}(Q(\cdot))$ as a linear operator, which facilitates subsequent calculations.
E.1 Eq. (1)
@@
$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\left({U_{A}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}\right)\right].$ (73)
@@
$\overline{Y}=\sum_{\sigma,\pi}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right).$ (75)
@@
$F_{1}(\sigma,E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right)$
@@
$\overline{Y}=\sum_{\sigma,\pi}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right)F_{1}(\sigma,E),$ (77)
@@
$\overline{Y}=\frac{4(D_{A}^{2}D_{E}^{2}-3D_{A}D_{E}-6D_{E}^{2}+6)}{D_{A}(D_{A}^{2}-9)D_{E}^{3}}.$ (78)
E.2 Eq. (4)
@@
$\overline{Y}=\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\right)\right].$ (81)
@@
$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right).$ (82)
The summation over $\vec{i},\vec{i^{\prime}}$ can be done similarly in Eq. (E.1):
$F_{2}(\sigma_{A},\sigma_{B},E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right)$
@@
$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{2}(\sigma_{A},\sigma_{B},E).$ (84)
@@
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\otimes{I_{C}}^{\otimes 4}\right)\right].$
@@
$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}\times$ (89)
@@
The summation over $\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}$ can be done in the similar way as those Eq. (E.1) and Eq. (E.2).
@@
::=:= $\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right)\times$
@@
$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g).$ (91)

### v3

@@ Appendix E Factorized unitary acting on three subregions
Figure 10: Numerical results for $\overline{Y^{\rm lin}}$ under brickwork circuits of varying depths and gate ranges, as depicted in Fig. 9.
Appendix F Magic injection with shallow-depth brickwork circuits
In this section, we replace the Haar random unitary gate acting on $A$ with a shallow-depth brickwork circuit. First, we still consider the initial state shown in Fig. 1(d) of the main text, followed by one layer of local gates acting on subregion $A$. Since the initial state factorizes into single-qubit states and Bell-pair states, the average $1-\overline{Y^{\rm lin}}$ also factorizes into a product of contributions from single-qubit and Bell states. Thus, it is easy to see that, $1-\overline{Y^{\rm lin}}\propto 2^{-E}$ still holds true in this case.
However, if $U_{A}$ is not Haar random, the specific choice of initial state in Fig. 1(d) is no longer the most generic one. To construct the most general stabilizer initial state with tunable entanglement, notice that since any bipartite stabilizer state is local-unitary equivalent to the state in Fig. 1(d), we can simply recover a generic stabilizer state by acting on Fig. 1(d) with random Clifford unitaries $U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}$. Since $U^{\rm Cliff}_{B}$ commutes with subsequent magic injection gates supported on $A$ and leaves $Y$ invariant, we can simply apply $U^{\rm Cliff}_{A}$. We then consider brickwork circuits of varying depths and gate ranges, as summarized in Fig. 9. We then numerically compute the average $\overline{Y^{\rm lin}}$ corresponding to each circuit architecture, and the results are shown in Fig. 10. We find that the qualitative picture of our analytical result is still valid for the brickwork circuit, in that the average $1-\overline{Y^{\rm lin}}$ decays exponentially with $E$. However, notice that for extremely shallow depths (single layer and two layers), the curve bends upward and saturates, which indicates magic injection that is below capacity. As the circuit depth increases, the behavior of $\overline{Y^{\rm lin}}$ approaches that of a Haar random $U_{A}$.
Appendix G Unitary stabilizer Rényi entropy
Here we develop the formalism for unitary magic and T count based on ideas similar to the stabilizer Rényi entropy and unitary nullity.
Definition G.1.
Let $U$ be an $N$-qubit unitary, then the $\alpha$-th unitary stabilizer Rényi entropy of $U$ is defined as
[…]
where $P_{i}\in P_{N}$.
For a shorthand notation, let $Q_{\alpha}:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes 2\alpha}$, then we can write it as:
$H_{\alpha}(U)=\frac{1}{1-\alpha}\text{log}[{\text{tr}}(Q_{\alpha}U^{\otimes 2\alpha}Q_{\alpha}{U^{\dagger\otimes 2\alpha}})]-2N\frac{1+\alpha}{1-\alpha}.$ (75)
As explained in the main text, the quantity $H_{\alpha}$ can be viewed as the $\alpha$-th Rényi entropy of the probability distribution associated with the weight $|c_{ij}|^{2}$ of $UP_{i}U^{\dagger}=\sum_{j}c_{ij}P_{j}$ over all $4^{N}$ Pauli strings.
G.1 The relationship between unitary SRE and state SRE
Consider a $2N$-qubit maximally entangled state between the original system and an auxiliary system, where a unitary $U$ acts on the $N$-qubit subsystem. The unitary stabilizer Rényi entropy $H_{\alpha}(U)$ is directly related to the state stabilizer Rényi entropy $M_{\alpha}$, defined as:
$M_{\alpha}(\rho):=\frac{1}{1-\alpha}\text{log}[\text{tr}({\rho}^{\otimes 2\alpha}Q_{\alpha})]-N\frac{1}{1-\alpha}.$ (76)
Theorem 3.
Suppose $U$ is an $N$-qubit unitary operator. Let $|U\rangle_{AB}=U_{A}\otimes I_{B}|{{\rm Bell}\rangle}^{\otimes N}_{AB})$ be the Choi state of $U$ where $|A|=|B|=2^{N}$ and $|\mathrm{Bell}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ is maximally entangled between $A$ and $B$ Then
$M_{\alpha}(|U\rangle)=H_{\alpha}(U).$
Proof.
Consider a $2N$-qubit maximally entangled state between the original system $A$ and an auxiliary system $B$ of equal size: $\rho_{0}=|{\rm Bell}\rangle\langle{\rm Bell}|^{\otimes N}$. Recall that a $2$-qubit Bell state can be written as:
$|{\rm Bell}\rangle\langle{\rm Bell}|=\frac{1}{4}(II+XX+ZZ-YY).$ (77)
[…]
$\rho_{0}=\frac{1}{2^{2N}}\sum_{P_{i}\in P_{N}}\underbrace{P_{i}}_{A}\underbrace{P_{i}}_{B}.$ (78)
Now we consider the stabilizer Rényi entropy of state $|U\rangle=U_{A}\otimes I_{B}\ |{\rm Bell}\rangle_{AB}^{\otimes N}$:
$M_{\alpha}(|U\rangle)=\frac{1}{1-\alpha}\text{log}[\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})]-2N\frac{1}{1-\alpha},$ (79)
the last term is not $N\frac{1}{1-\alpha}$ because $|U\rangle$ is a $2N$-qubit state. The key problem is to calculate the quantity inside the logarithm:
$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (80)
$=$ $\sum_{P_{i_{1}},\dots,P_{i_{2\alpha}}\in P_{N}}2^{-4\alpha N}\text{tr}(Q^{AB}_{\alpha}{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{A}\otimes\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{B}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}).$
From the definition of $Q_{\alpha}$, $Q^{AB}_{\alpha}$ can be written as a tensor product $Q^{AB}_{\alpha}=Q^{A}_{\alpha}\otimes Q^{B}_{\alpha}$, then Eq. (80) can be written as:
$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (81)
$=$ $\sum_{P_{i_{1}},\dots,P_{i_{2\alpha}}\in P_{N}}2^{-4\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{A}{U_{A}}^{\dagger\otimes 2\alpha})\text{tr}(Q_{\alpha}{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{B}{I_{B}}^{\otimes 2\alpha}),$
because $Q_{\alpha}:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes{2\alpha}}$, it will select those $P_{i_{1}}\cdots P_{i_{2\alpha}}$ with $P_{i_{1}}=P_{i_{2}}=\cdots=P_{i_{2\alpha}}$. Then Eq. (81) further simplifies:
$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (82)
$=$ $\sum_{P_{i}\in P_{N}}2^{-4\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}\underbrace{P_{i}\cdots P_{i}}_{2\alpha}{U_{A}}^{\dagger\otimes 2\alpha})\text{tr}(Q_{\alpha}{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i}\cdots P_{i}}_{2\alpha}{I_{B}}^{\otimes 2\alpha})$
$=$ $2^{-2\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha}).$
Then $M_{\alpha}(|U\rangle)$ can be written as:
$M_{\alpha}(|U\rangle)=$ $\frac{1}{1-\alpha}\text{log}[\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})]-2N\frac{1}{1-\alpha}$ (83)
$=$ $\frac{1}{1-\alpha}\text{log}[2^{-2\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha)}]-2N\frac{1}{1-\alpha}$
$=$ $\frac{1}{1-\alpha}\text{log}[\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha)}]-2N\frac{1+\alpha}{1-\alpha},$
this is exactly the definition of $H_{\alpha}(U)$ in Eq. (75), so we get $H_{\alpha}(U)=M_{\alpha}(|U\rangle)$. ∎
We have thus established that the unitary stabilizer Rényi entropy $H_{\alpha}(U)$ is equal to the state Rényi stabilizer entropy of a maximally entangled state between the original system and an auxiliary system of equal size, acted upon by $U$. Notice that this result does not involve averaging over a particular unitary ensemble, and holds for each and arbitrary $U$.
G.2 Lower bound on the unitary nullity and $T$ count
In this section, we establish the full inequality of Eq. (7) of the main text. Let $p_{ij}(U):={\text{tr}}^{2}(P_{i}UP_{j}U^{\dagger})/2^{4N}$. Because $\sum_{i,j}p_{ij}=1$, the coefficients $p_{ij}$ form a normalized probability distribution. Then the $\alpha$-th unitary stabilizer Rényi entropy we have defined can be written in the following way:
$H_{\alpha}(U):=\frac{1}{1-\alpha}\text{log}\sum_{i,j}{p_{ij}(U)}^{\alpha}-2N.$ (84)
To connect the unitary SRE to unitary nullity, we first note the following lemma.
Theorem 4.
Let $t(U)$ be the T-count of $U$, i.e., the number of T gates needed to synthesize $U$ in a Clifford+T decomposition and let $\nu(U)$ be the unitary nullity. Then $H_{0}(U)\leq\nu(U)\leq t(U)$.
Proof.
Define $s(U):=\{P_{i}|\ UP_{i}U^{\dagger}\in P_{N}\}$ and $g(U):=Us(U)U^{\dagger}$. Both $s(U)$ and $g(U)$ are subgroups of $P_{N}$. Recall the definition of unitary nullity [51] is $v(U):=2N-\text{log}|s(U)|$. We will show that, for a fixed $P_{i}$, the number of non-zero $p_{ij}$ is upper bounded by $\frac{2^{2N}}{|s(U)|}$.
The group $g(U)$ acts on the operator space via conjugation, $R(X):=RXR,\ R\in g(U)$. For any $R\in g(U)$, there exists $Q\in s(U)$ with $R=UQU^{\dagger}$. Thus for any Pauli string $P_{i}$ and any $R\in g(U)$, we have:
$R(UP_{i}U^{\dagger})=RUP_{i}U^{\dagger}R=U(QP_{i}Q)U^{\dagger},$ (85)
for some $Q\in s(U)$. Since Pauli strings either commute or anticommute, that is, $QP_{i}=\pm P_{i}Q$, we must have
$R(UP_{i}U^{\dagger})=\pm UP_{i}U^{\dagger}.$ (86)
Thus, regarding $UP_{i}U^{\dagger}$ as a vector in operator space and $R(\cdot)$ as a linear transformation, we see that $UP_{i}U^{\dagger}$ is a common eigenvector of all the $R\in g(U)$, with eigenvalues $\pm 1$. Expanding $UP_{i}U^{\dagger}$ in Pauli basis,
$UP_{i}U^{\dagger}=\sum_{S\in P_{N}}c_{S}S,$ (87)
and applying $R(\cdot)$ gives
$R(UP_{i}U^{\dagger})=\sum_{S}c_{S}(RSR)=\pm\sum_{S}c_{S}S.$ (88)
Since $RSR=\pm S$, Eq. (86) imposes the constraint that all Pauli strings in the above expansion with $c_{S}\neq 0$ must share the same commutation relation with $R$. Modulo a global phase, the space of $P_{N}$ is isomorphic to the binary symplectic space $\mathbb{F}_{2}^{2N}$. Each Pauli string $S=X_{1}^{s_{1}}Z_{1}^{s_{2}}X_{2}^{s_{3}}Z_{2}^{s_{4}}...X_{N}^{s_{2N-1}}Z_{N}^{s_{2N}}$ corresponds to a vector $s=(s_{1},s_{2},...,s_{2N})$, and each $R\in g(U)$ can be equivalently represented as a vector $r\in\mathbb{F}_{2}^{2N}$. The commutation relation between $R$ and $S$ is encoded by the symplectic form:
$\varepsilon(R,S)=(-1)^{\langle r,s\rangle},\quad\langle r,s\rangle:=\sum^{N}_{i=1}r_{2i-1}s_{2i}+r_{2i}s_{2i-1}\quad({\rm mod}\ 2)$ (89)
where the $\langle r,s\rangle$ is the symplectic inner product of $r,s$. Therefore, the constraint that all nonzero $c_{S}$ share the same commutation relation with $R\in g(U)$ becomes a system of linear equations:
$\langle r,s\rangle=b(r)\in\{0,1\}\quad\forall\ r\in G,$ (90)
where $G\subseteq\mathbb{F}^{2N}_{2}$ is the vector space corresponding to $g(U)$, with $\text{dim}G=\text{log}|g(U)|=\text{log}|s(U)|$. Thus the solution space lies within the affine translation of the symplectic orthogonal complement $G^{\perp}$, whose size is $|G^{\perp}|=2^{2N-\text{dim}G}=\frac{2^{2N}}{|s(U)|}$. Hence, the number of nonzero coefficients $c_{S}$ is at most $\frac{2^{2N}}{|s(U)|}$. Summing over all Pauli strings $P_{i}$, we obtain:
$H_{0}(U)=\text{log}(\#\{p_{ij}\neq 0\})-2N\leq\text{log}(\frac{2^{4N}}{|s(U)|})-2N=v(U).$ (91)
Since $t(U)\geq v(U)$ [51], it follows that $t(U)\geq v(U)\geq H_{0}(U)$. ∎
Corollary G.1.
For any unitary $U$, its $T$-count is lower bounded such that $t(U)\geq H_{\alpha}(U)$.
Proof.
Recall that Rényi entropy is non-increasing in $\alpha$ such that $H_{\infty}(U)\leq H_{\alpha}(U)\leq H_{0}(U)$ where $H_{0}(U)=\text{log}(\#\{p_{ij}\neq 0\})-2N$. ∎
Although $H_{\alpha}(U)$ is not as tight a lower bound compare to unitary nullity, the calculation of $H_{\alpha}(U)$ does not involve any optimization. Furthermore, if $U_{A}$ has a finer structure, such as a finite depth circuit, then the stabilizer Renyi entropy of its Choi state can be computed efficiently using tensor network methods. Hence we obtain a computable measure of the unitary magic and a computable lower bound of $t(U)$.
Appendix H Exact results
@@
In this section, we present an alternative analytical method to derive exact expressions of $\overline{Y^{\rm lin}}$ in all cases discussed in the main text [^2] We acknowledge Alioscia Hamma for providing insightful notes, which served as the primary inspiration for the analytical method presented in this section.. Recall the linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y^{\rm lin}(\rho):=1-\frac{1}{D}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$. where $D:={\rm dim}\ \mathcal{H}$. This quantity admits an equivalent expression:
$Y^{\rm lin}(\rho)=1-\frac{1}{D}{\rm tr}(Q\rho^{\otimes 4}),$ (92)
@@
where $Q:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes 4}$. For a Hilbert space $\mathcal{H}=\otimes^{n}_{i=1}\mathcal{H}_{i}$, the operator $Q$ factories as $Q=\otimes^{n}_{i=1}Q_{i}$. We can then interpret the map ${\rm tr}(Q(\cdot))$ as a linear operator, which facilitates subsequent calculations.
H.1 Eq. (2)
@@
$\overline{Y^{\rm lin}}=\mathbb{E}_{U_{A}}\ Y^{\rm lin}\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\left({U_{A}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}\right)\right].$ (95)
@@
$1-\overline{Y^{\rm lin}}=\sum_{\sigma,\pi}\sum_{\vec{i},\vec{i^{\prime}}}\frac{1}{D}2^{-4E}V^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right).$ (97)
@@
$F_{1}(\sigma,E)$ $:=$ $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right)$
@@
$1-\overline{Y^{\rm lin}}=\sum_{\sigma,\pi}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right)F_{1}(\sigma,E),$ (99)
@@
$\overline{Y^{\rm lin}}=1-\frac{4(D_{A}^{2}D_{E}^{2}-3D_{A}D_{E}-6D_{E}^{2}+6)}{D_{A}(D_{A}^{2}-9)D_{E}^{3}}.$ (100)
H.2 Eq. (LABEL:eq:UaUb)
@@
$\overline{Y^{\rm lin}}=\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)=1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\right)\right].$ (103)
@@
$1-\overline{Y^{\rm lin}}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}\frac{1}{D}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right).$ (104)
The summation over $\vec{i},\vec{i^{\prime}}$ can be done similarly in Eq. (H.1):
$F_{2}(\sigma_{A},\sigma_{B},E)$ $:=$ $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right)$
@@
$1-\overline{Y^{\rm lin}}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}2^{-4E}\frac{1}{D}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{2}(\sigma_{A},\sigma_{B},E).$ (106)
@@
$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\otimes{I_{C}}^{\otimes 4}\right)\right].$
@@
$1-\overline{Y^{\rm lin}}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}\frac{1}{D}\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}\times$ (111)
@@
The summation over $\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}$ can be done in the similar way as those Eq. (H.1) and Eq. (H.2).
@@
$:=$ $\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right)\times$
@@
$1-\overline{Y^{\rm lin}}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}\frac{1}{D}\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g).$ (113)